# Deep Linear Probe Generators for Weight Space Learning

### v1

@@ 1 Introduction
The growing importance and popularity of neural networks has led to the development of several model hubs (e.g. HuggingFace, CivitAI), where more than a million models are now publicly available. Treating them as a new data modality presents new opportunities for machine learning. Specifically, as not all neural networks include information about their training, developing methods to automatically learn from weights is becoming important. For instance, given an undocumented model, it is interesting to know its generalization error [38] or its training dataset [10]. While some of these questions could be answered by evaluating the model on many labelled samples, this is often impractical as the data may be unavailable or unknown. Here, we want to answer these questions without access to the models true data distribution and under minimal computational effort.
Learning from weights is essentially similar to the well studied problem of binary code analysis [36], where the goal is to predict the function of an unknown software using its binary code. In both cases, the task is to understand an unknown complex function specified by many parameters. Binary code analysis approaches generally fall into two categories: static and dynamic. Static methods [36] aim to understand a function without running the binary code. Dynamic code analysis [11, 4] runs the code on inputs provided by the user and analyzes its outputs to understand what the code does. Similarly, in weight space learning, there are two main types of methods: mechanistic approaches [38, 26, 41, 25] aim to understand model weights without running the model, while probing approaches [17, 21] represent models by their responses to a set of well selected inputs. Despite dynamic methods often performing better for binary code analysis, in the context of weight space learning, probing is still under utilized.
@@
To this end, we propose Deep Linear Probe Generators (ProbeGen) as a simple and effective solution. ProbeGen factorizes its probes into two parts, a per-probe latent code and a global probe generator. The generator offers two key benefits: (i) It helps sharing information across multiple probes, and (ii) can implicitly introduce an inductive bias into the probes. For example, in images, hierarchical and convolutional layers create a local structure. Finally, by observing the learned probes we hypothesize they are not necessarily semantic, and owe some of their expression to low-level structures. We then find that the non-linear activation functions, which increase expressivity, actually degrade the learned probes. Our final approach therefore consists of a deep linear network [1], with data-dependent biases. Our linear generators produce probes that achieve state-of-the-art performance on common weight space learning tasks.
@@ 2 Related Work
A recent line of works have focused on training models to process the weights of diverse model populations (known as Model Zoos, e.g., [33]) to predict undocumented properties of a model. These properties include classifying the training dataset or predicting the generalization error. Mechanistic methods [25, 26, 34, 12, 38] represent models using the mechanics of their inner workings. One approach [7, 34, 32] is to learn standard architectures over raw weights. However, weights exhibit permutation symmetries between neurons [26] which these architectures do not explicitly account for, although some these methods use augmentations [34, 32] to encourage permutation invariance. Another approach [38, 10, 31] maps the weights to a low dimensional embedding by a set of weight statistics, which are completely invariant to permutations, but have limited expressivity as they ignore the inner relations between neurons. Recently, a line of works [26, 41, 42, 21, 25] focus on specially designing permutation equivariant architectures for processing neural networks. A dominant approach uses graph based architectures [21, 25] modeling a neural network as a computational graph, where every neuron is a node. They then train Graph Neural Network (GNN) [14, 20] or Transformer [39, 8] modules, which are equivariant by design, to analyze the computational graph. Most recently, some equivariant approaches [21, 17] also included learned probes. In this work we take a deep look into probing methods, and their failure points.
Another research thrust developed new applications for weight space learning. Some works [16, 2, 27] encode the parameters of neural networks, mainly for generating, modifying or compressing weights matrices. Others [13, 9, 35] use weights for advanced image generation capabilities. More recently, Carlini et al. [6] proposed recovering the exact black-boxed weights of a neural network layer, and Horwitz et al. [18] demonstrated recovering entire pre-trained models when the weights of multiple fine-tuned models of the same Model Tree [19] are available.
@@ 3 Background
The Challenge. While a naive solution would be to apply a standard neural architecture directly to the weights, this idea encounters serious setbacks. The dimension of the weight vector is very high, but more fundamentally, the ordering of the neurons in each layer is arbitrary. Hence, different permutations of the neurons result in functionally identical models. Standard architectures such as MLPs and CNNs do not respect such symmetries. One popular approach by Unterthiner et al. [38] computes permutation-invariant statistics for the flattened weights and biases of each layer, then trains a standard classifier on them. These statistics lose much of the information contained in the weights, limiting the potential of this approach. Another direction is learning with equivariant architectures [26, 41] which respect the permutation invariance, e.g., graph neural networks [21, 25]. However, these methods [21] treat each neuron of the model as a token, and scaling them to large architectures is challenging (see Sec. 5.1).
@@
Assuming that we want to predict an attribute $y$ of a network $f$, probing methods [17, 21] optimize a set of $k$ probes $(p_{1},...,p_{k})$, and feed them into the network. They then train a classifier $C$ on the concatenation of the outputs. The prediction $\hat{y}$ is:
@@
Probing methods learn the parameters of each probe $p$ directly by latent optimization [5]. Each probe provides some information about the model attributes, and learning diverse and discriminative probes is key for obtaining a useful representation. The classifier $C$ leverages information from all probes, and is typically trained by cross-entropy for classification and mean squared error for regression.
@@ 4 Weight Space Learning with Deep Linear Probe Generators
Our initial hypothesis is that probing methods, when done right, hold significant potential. Much like binary code files, neural networks are unknown and highly complex functions. Drawing inspiration from binary code analysis, where dynamic approaches [11, 4] are more common than static ones [36], we believe that running neural networks, i.e., probing, is a promising approach for weight space learning. We begin with $2$ preliminary experiments to test the quality and potential of probing approaches.
@@ 4.1 A Simple Probing Baseline
As we believe probing should be an effective way of analyzing neural networks, we begin by testing its raw capabilities. We take a vanilla probing approach, without any enhancements or modifications, that optimizes all probes $p_{1},p_{2},...,p_{k}$ with latent optimization (i.e., optimizing their values directly) and uses a simple MLP classifier for $C$. We compare this vanilla probing to its top competitors. First, a graph based approach (Neural Graphs) [21] which treats each neuron of the network as a node and operates a transformer on the resulting computational graph. Second, weight statistics (StatNN) [38] which extracts simple statistics from the flattened weights and biases of a network and trains a simple predictor over them. We test all approaches on $4$ popular benchmarks. For dataset classification, we measure the accuracy for MNIST [23] INRs digit prediction and Fashion-MNIST [40] INRs class prediction, both provided by Navon et al. [26]. For generalization error prediction, we measure the Kendall’s $\tau$ for CIFAR10-GS [38] and CIFAR10-Wild-Park [21]. The results are presented in Tab. 1. It is clear that (i) with enough probes, vanilla probing is able to perform better than a graph based approach that does not use probing. (ii) Graph based approaches become comparable to vanilla probing only when incorporating probing features. This shows the promise of probing methods. Additionally, in Sec. 5.1 we demonstrate that probing also requires much less computational resources than graph based methods.
Table 1: Simple Probing vs. Other Approaches. We compare a simple probing approach to previous graph based and mechanistic approaches. We average the results over $5$ different seeds. For probing, we experiment with different numbers of probes (in brackets).
@@
StatNN (0) 0.398 ±$\pm 0.001$ 0.418 ±$\pm 0.002$ 0.914 ±$\pm 0.000$ 0.719 ±$\pm 0.010$
Neural Graphs (0) 0.923 ±$\pm 0.003$ 0.727 ±$\pm 0.006$ 0.935 ±$\pm 0.000$ 0.817 ±$\pm 0.007$
Neural Graphs (64) 0.967 ±$\pm 0.002$ 0.736 ±$\pm 0.012$ 0.938 ±$\pm 0.001$ 0.888 ±$\pm 0.009$
Neural Graphs (128) 0.976 ±$\pm 0.001$ 0.745 ±$\pm 0.008$ 0.938 ±$\pm 0.000$ 0.885 ±$\pm 0.005$
Vanilla Probing (64) 0.873 ±$\pm 0.026$ 0.784 ±$\pm 0.017$ 0.933 ±$\pm 0.001$ 0.885 ±$\pm 0.008$
Vanilla Probing (128) 0.955 ±$\pm 0.005$ 0.808 ±$\pm 0.006$ 0.936 ±$\pm 0.001$ 0.889 ±$\pm 0.008$
@@
Having shown the merit in probing approaches, it is interesting to understand the quality of the probes themselves. To do so, we replace the learned probes by a set of unlearned inputs, training only the classifier $C$. We assume no knowledge about the training data, by testing random synthetic data probes. We fix the number of probes $k$ in all cases. In the MNIST and FMNIST INRs tasks, we choose points from a uniform distribution between $[[-1,1],[-1,1]]$. In the CIFAR10 cases, we select $k$ random synthetic images from the Dead Leaves [3, 24] dataset. As seen in Fig. 2, these images are synthetic, unlearned and highly dissimilar to real images. We choose these images as they include some structure yet are still far from being realistic. The results are presented in Tab. 2. It is clear that random probes are comparable to learned ones. We conclude that current probing techniques find suboptimal probes. To understand why these learned probes perform worse, we first observe them in Fig. 3(a). The probes show low-level, almost adversarial patterns, which can be highly expressive [15]. We therefore hypothesize vanilla probing tends to overfit. In Sec. 5.2 we demonstrate that our final method indeed overfits less than vanilla probing, and that more expressive methods hurt performance in these tasks.
@@ 4.3 Deep Linear Probe Generators
Table 2: Learned vs. Out of Distribution Probes. Comparison learned probes and probes from randomly selected data. We average the results over $5$ different seeds.

### v2

@@ 1 Introduction
The growing importance and popularity of neural networks has led to the development of several model hubs (e.g. HuggingFace, CivitAI), where more than a million models are now publicly available. Treating them as a new data modality presents new opportunities for machine learning. Specifically, as not all neural networks include information about their training, developing methods to automatically learn from weights is becoming important. For instance, given an undocumented model, it is interesting to know its generalization error (Unterthiner et al., 2020) or its training dataset (Dupont et al., 2022). While some of these questions could be answered by evaluating the model on many labelled samples, this is often impractical as the data may be unavailable or unknown. Here, we want to answer these questions without access to the models true data distribution and under minimal computational effort.
Learning from weights is essentially similar to the well studied problem of binary code analysis Shoshitaishvili et al. (2016), where the goal is to predict the function of an unknown software using its binary code. In both cases, the task is to understand an unknown complex function specified by many parameters. Binary code analysis approaches generally fall into two categories: static and dynamic. Static methods Shoshitaishvili et al. (2016) aim to understand a function without running the binary code. Dynamic code analysis Egele et al. (2008); Bayer et al. (2006) runs the code on inputs provided by the user and analyzes its outputs to understand what the code does. Similarly, in weight space learning, there are two main types of methods: mechanistic approaches Unterthiner et al. (2020); Navon et al. (2023a); Zhou et al. (2024a); Lim et al. (2023); Kalogeropoulos et al. (2024) aim to understand model weights without running the model, while probing approaches Herrmann et al. (2024); Kofinas et al. (2024) represent models by their responses to a set of well selected inputs. Despite dynamic methods often performing better for binary code analysis, in the context of weight space learning, probing is still under utilized.
@@
To this end, we propose Deep Linear Probe Generators (ProbeGen) as a simple and effective solution. ProbeGen factorizes its probes into two parts, a per-probe latent code and a global probe generator. The generator offers two key benefits: (i) It helps sharing information across multiple probes, and (ii) can implicitly introduce an inductive bias into the probes. For example, in images, hierarchical and convolutional layers create a local structure. Finally, by observing the learned probes we hypothesize they are not necessarily semantic, and owe some of their expression to low-level structures. We then find that the non-linear activation functions, which increase expressivity, actually degrade the learned probes. Our final approach therefore consists of a deep linear network (Arora et al., 2019), with data-dependent biases. Our linear generators produce probes that achieve state-of-the-art performance on common weight space learning tasks.
@@ 2 Related Work
Weight Space Learning.
A recent line of works have focused on training models to process the weights of diverse model populations (known as Model Zoos, e.g., (Schürholt et al., 2022)) to predict undocumented properties of a model. These properties include classifying the training dataset (e.g., the class of an image used to train an Implicit Neural Representation) or predicting the generalization error. Mechanistic methods (Lim et al., 2023; Navon et al., 2023a; Schürholt et al., 2024; Eilertsen et al., 2020; Unterthiner et al., 2020) represent models using the mechanics of their inner workings. One approach (De Luigi et al., 2023; Schürholt et al., 2024; 2021) is to learn standard architectures over raw weights. However, weights exhibit permutation symmetries between neurons (Navon et al., 2023a) which these architectures do not explicitly account for, although some of these methods use augmentations (Schürholt et al., 2024; 2021) to encourage permutation invariance. Other methods proposed specialized augmentations (Shamsian et al., 2024) which can be applied to aligned networks (Shamsian et al., 2024; Navon et al., 2023b; Peña et al., 2023; Ainsworth et al., 2022) or to already aligned weight spaces (Lim et al., 2024b; a). Another approach (Unterthiner et al., 2020; Dupont et al., 2022; Salama et al., 2024) maps the weights to a low dimensional embedding by a set of weight statistics, which are completely invariant to permutations, but have limited expressivity as they ignore the inner relations between neurons. Recently, a line of works (Navon et al., 2023a; Zhou et al., 2024a; b; Kofinas et al., 2024; Lim et al., 2023; Tran et al., 2024; Kalogeropoulos et al., 2024) focus on specially designing permutation equivariant architectures for processing neural networks. A dominant approach uses graph based architectures (Kofinas et al., 2024; Lim et al., 2023; Kalogeropoulos et al., 2024) modeling a neural network as a computational graph, where every neuron is a node. They then train Graph Neural Network (GNN) (Gilmer et al., 2017; Kipf & Welling, 2016) or Transformer (Vaswani, 2017; Diao & Loynd, 2022) modules, which are equivariant by design, to analyze the computational graph. Most recently, some equivariant approaches (Kofinas et al., 2024; Herrmann et al., 2024) also included learned probes. In this work we take a deep look into probing methods, and their failure points.
Another research thrust developed new applications for weight space learning. Some works (Ha et al., 2016; Ashkenazi et al., 2022; Peebles et al., 2022) encode the parameters of neural networks, mainly for generating, modifying or compressing weights matrices. Others (Erkoç et al., 2023; Dravid et al., 2024; Shah et al., 2023) use weights for advanced image generation capabilities. More recently, Carlini et al. (2024) proposed recovering the exact black-boxed weights of a neural network layer, and Horwitz et al. (2024a) demonstrated recovering entire pre-trained models when the weights of multiple fine-tuned models of the same Model Tree Horwitz et al. (2024b) are available.
Implicit Neural Representations (INRs).
In recent years, INRs (Sitzmann et al., 2020; Tancik et al., 2020) have emerged as a new paradigm for representing data points with neural networks. These networks are used in various data modalities such as images (Ha, 2016), 3D shapes (Mescheder et al., 2019; Chen & Zhang, 2019), 3D scenes (Mildenhall et al., 2021), video (Li et al., 2021), audio (Sitzmann et al., 2020), etc. Specifically, in the context of images, INRs are neural networks trained on a single image, which given a pixel location $(x,y)$ return the value of that pixel in the training image. In this work, we classify the class of a training image an INR network was trained on, using only the INR’s weights.
@@ 3 Background
The Challenge. While a naive solution would be to apply a standard neural architecture directly to the weights, this idea encounters serious setbacks. The dimension of the weight vector is very high, but more fundamentally, the ordering of the neurons in each layer is arbitrary. Hence, different permutations of the neurons result in functionally identical models. Standard architectures such as MLPs and CNNs do not respect such symmetries. One popular approach by Unterthiner et al. (2020) computes permutation-invariant statistics for the flattened weights and biases of each layer, then trains a standard classifier on them. These statistics lose much of the information contained in the weights, limiting the potential of this approach. Another direction is learning with equivariant architectures (Navon et al., 2023a; Zhou et al., 2024a) which respect the permutation invariance, e.g., graph neural networks Kofinas et al. (2024); Lim et al. (2023). However, these methods (Kofinas et al., 2024) treat each neuron of the model as a token, and scaling them to large architectures is challenging (see Sec. 5.1).
@@
Assuming that we want to predict an attribute $y$ of a network $f$, probing methods (Herrmann et al., 2024; Kofinas et al., 2024) optimize a set of $k$ probes $(p_{1},...,p_{k})$, and feed them into the network. They then train a classifier $C$ on the concatenation of the outputs. The prediction $\hat{y}$ is:
@@
Probing methods learn the parameters of each probe $p$ directly by latent optimization (Bojanowski et al., 2017). Each probe provides some information about the model attributes, and learning diverse and discriminative probes is key for obtaining a useful representation. The classifier $C$ leverages information from all probes, and is typically trained by cross-entropy for classification and mean squared error for regression.
@@ 4 Weight Space Learning with Deep Linear Probe Generators
Our initial hypothesis is that probing methods, when done right, hold significant potential. Much like binary code files, neural networks are unknown and highly complex functions. Drawing inspiration from binary code analysis, where dynamic approaches Egele et al. (2008); Bayer et al. (2006) are more common than static ones Shoshitaishvili et al. (2016), we believe that running neural networks, i.e., probing, is a promising approach for weight space learning. We begin with $2$ preliminary experiments to test the quality and potential of probing approaches.
@@ 4.1 A Simple Probing Baseline
As we believe probing should be an effective way of analyzing neural networks, we begin by testing its raw capabilities. We take a vanilla probing approach, without any enhancements or modifications, that optimizes all probes $p_{1},p_{2},...,p_{k}$ with latent optimization (i.e., optimizing their values directly) and uses a simple MLP classifier for $C$. We compare this vanilla probing to its top competitors. First, a graph based approach (Neural Graphs) (Kofinas et al., 2024) which treats each neuron of the network as a node and operates a transformer on the resulting computational graph. Second, weight statistics (StatNN) (Unterthiner et al., 2020) which extracts simple statistics from the flattened weights and biases of a network and trains a simple predictor over them. We test all approaches on $4$ popular benchmarks. For dataset classification, we measure the accuracy for MNIST LeCun et al. (1998) INRs digit prediction and Fashion-MNIST Xiao et al. (2017) INRs class prediction, both provided by Navon et al. (2023a). For generalization error prediction, we use the CIFAR10-GS (Unterthiner et al., 2020) and CIFAR10-Wild-Park (Kofinas et al., 2024) benchmarks, measuring the Kendall’s $\tau$ metric as common in weight-space learning evaluation. The Kendall’s $\tau$ is a statistic used to measure the correlation agreement between two rankings, where $1$ indicates perfect correlation, and 1-1 indicates perfect negative correlation. The results are presented in Tab. 1. It is clear that (i) with enough probes, vanilla probing is able to perform better than a graph based approach that does not use probing. (ii) Graph based approaches become comparable to vanilla probing only when incorporating probing features. This shows the promise of probing methods. Additionally, in Sec. 5.1 we demonstrate that probing also requires much less computational resources than graph based methods.
Table 1: Simple Probing vs. Other Approaches. We compare a simple probing approach to previous graph based and mechanistic approaches. We average the results over 55 different seeds. For probing, we experiment with different numbers of probes (in brackets).
@@
StatNN (0) 0.398 $\pm 0.001$ 0.418 $\pm 0.002$ 0.914 $\pm 0.000$ 0.719 $\pm 0.010$
Neural Graphs (0) 0.923 $\pm 0.003$ 0.727 $\pm 0.006$ 0.935 $\pm 0.000$ 0.817 $\pm 0.007$
Neural Graphs (64) 0.967 $\pm 0.002$ 0.736 $\pm 0.012$ 0.938 $\pm 0.001$ 0.888 $\pm 0.009$
Neural Graphs (128) 0.976 $\pm 0.001$ 0.745 $\pm 0.008$ 0.938 $\pm 0.000$ 0.885 $\pm 0.005$
Vanilla Probing (64) 0.873 $\pm 0.026$ 0.784 $\pm 0.017$ 0.933 $\pm 0.001$ 0.885 $\pm 0.008$
Vanilla Probing (128) 0.955 $\pm 0.005$ 0.808 $\pm 0.006$ 0.936 $\pm 0.001$ 0.889 $\pm 0.008$
@@
Having shown the merit in probing approaches, it is interesting to understand the quality of the probes themselves. To do so, we replace the learned probes by a set of unlearned inputs, training only the classifier $C$. We test both: (i) probes with no knowledge about the training data, selected from random synthetic data, and (ii) probes from in-distribution data, selected from the training set of the networks. We fix the number of probes $k$ in all cases. In the MNIST and FMNIST INRs tasks, synthetic and in-distribution probes are the same, where we choose points from a uniform distribution between [[1,1],[1,1]][[-1,1],[-1,1]]. In the CIFAR10 cases, we select $k$ random synthetic images from the Dead Leaves (Baradad Jurjo et al., 2021; Lee et al., 2001) dataset as synthetic probes, and $k$ random images from CIFAR10 as in-distribution probes. As seen in Fig. 2, Dead Leaves images are synthetic, unlearned and highly dissimilar to real images. We therefore choose these images as our synthetic probes as they include some structure, yet are still far from being realistic. The results are presented in Tab. 2. It is clear that random probes are comparable to learned ones. We conclude that current probing techniques find suboptimal probes. To understand why these learned probes perform worse, we first observe them in Fig. 3(a). The probes show low-level, almost adversarial patterns, which can be highly expressive (Goodfellow et al., 2014). We therefore hypothesize vanilla probing tends to overfit. In Sec. 5.2 we demonstrate that our final method indeed overfits less than vanilla probing, and that more expressive methods hurt performance in these tasks.
Table 2: Learned vs. Out of Distribution Probes. Comparison of learned probes, in-distribution data probes and probes from randomly selected data. We average the results over 55 different seeds.
Accuracy Kendall’s $\tau$ ($\uparrow$)
# Probes Method MNIST FMNIST CIFAR10-GS CIFAR10 Wild Park
64 Learned Probes 0.873 $\pm 0.026$ 0.784 $\pm 0.017$ 0.933 $\pm 0.001$ 0.885 $\pm 0.008$
Synthetic Probes 0.899 $\pm 0.022$ 0.832 $\pm 0.010$ 0.918 $\pm 0.001$ 0.882 $\pm 0.007$
In-Dist. Probes 0.899 $\pm 0.022$ 0.832 $\pm 0.010$ 0.939 $\pm 0.000$ 0.922 $\pm 0.008$
128 Learned Probes 0.955 $\pm 0.005$ 0.808 $\pm 0.006$ 0.936 $\pm 0.001$ 0.889 $\pm 0.008$
Synthetic Probes 0.959 $\pm 0.006$ 0.856 $\pm 0.007$ 0.917 $\pm 0.001$ 0.893 $\pm 0.014$
In-Dist. Probes 0.959 $\pm 0.006$ 0.856 $\pm 0.007$ 0.941 $\pm 0.001$ 0.930 $\pm 0.011$
(a) Latent Optimization
(b) ProbeGen
Figure 3: Learned Probes. Optimized probes by (a) Latent Optimization and (b) ProbeGen for the CIFAR10 Wild Park benchmark. We show the same probes for identical runs (including the seed), except the generator module.
@@ 4.3 Deep Linear Probe Generators
Deep Linear Networks.
Non-linear activations such as ReLU are the engine that makes deep networks very expressive. However, in this case, we would like to add regularization (see Sec. 4.2) rather than increasing expressivity. We therefore remove the activations from our generator, keeping only stacked linear layers. Work by Arora et al. (2019) showed that when using SGD, deep linear networks have an implicit regularization effect. We therefore use deep linear networks in our approach, i.e., stacked linear layers but without the non-linear activations between them. In Sec. 5.2, we show that removing the activations reduces overfitting and therefore also enhances performance.
Data-Specfic Inductive Bias.
In the case that our target model $f$ takes structured inputs such as images, we hypothesize that presenting probes from a similar distribution will achieve higher accuracy and reduce overfitting. As we do not know the data distribution, we cannot learn an accurate generative model for it. Still, we can introduce data-specific inductive biases into the generator. For example, we present a generator for images, where we simply stack 2D convolutional layers, similarly to DCGAN (Radford et al., 2015). While this does not guarantee natural image statistics, it at least encourages some local structure and multi-scale hierarchy, both of which are some of the most important image characteristics.
5 Experiments
Our evaluation follows the standard protocol for weight space learning. We evaluate on two tasks: (i) CNNs generalization error prediction and (ii) detecting the training classes of images based on INR networks trained on them. We include experiments on small-scale established benchmarks as well as a new larger-scale Model Zoo which we present, using ResNet18(He et al., 2016) models.
Baselines.
We compare our method, ProbeGen, with StatNN (Unterthiner et al., 2020), DWS(Navon et al., 2023a), NFN (Zhou et al., 2024a), ScaleGMN (Kalogeropoulos et al., 2024) and Neural Graphs (Kofinas et al., 2024). StatNN computes $7$ statistics for the weights and biases of each layer, concatenates them and trains a gradient boosted tree method on this representation. DWS (Navon et al., 2023a) and NFN (Zhou et al., 2024a) train linear, permutation equivariant layers with non-linearities between them, but DWS is not applicable for CNNs. ScaleGMN (Kalogeropoulos et al., 2024) treats neural networks as computational graphs, accounting for permutation symmetries and scaling symmetries. Neural Graphs (Kofinas et al., 2024) is another graph-based approach, where each bias is a node and the weights are the matching edges between these nodes. This method then trains a transformer on the created graph, so the attention score between a pair of neurons depends on the weight that connects them. While Neural-Graphs do not account for scaling symmetries like the ScaleGMN baseline, it enriches each neuron’s representation by using probing features.
Datasets.
We evaluate on $4$ established datasets. For training data prediction we choose the MNIST and FMNIST implicit neural representation (INR) benchmarks (Navon et al., 2023a). Both datasets were formed by training an INR Sitzmann et al. (2020) model for each image of the original dataset. The goal is predicting the class of an image given its INR network. For generalization error prediction, we used the CIFAR10-GS (Unterthiner et al., 2020) and CIFAR10 Wild Park tasks Kofinas et al. (2024). These datasets consists of thousands of small CNN models (353-5 layers), each trained separately on CIFAR10 (Krizhevsky et al., 2009). We use accuracy for weight classification and Kendall’s $\tau$ for regression.
Metrics.
For INR class prediction, we use simple accuracy to measure performance. For predicting the generalization error of CNNs, we follow standard evaluation protocol and use the Kendall’s $\tau$ metric, which measures the agreement between two rankings. Similar to Pearson’s correlation, with Kendall’s $\tau$, values near $1$ indicate strong positive correlation, values near 1-1 indicate strong negative correlation, and values near 0 indicate no correlation.
Table 3: Results for Small Scale Benchmarks. Comparison of ProbeGen, to graph based, mechanistic approaches and latent optimized probes. We average the results over 55 different seeds.